If cosx/cosy=m/n prove that mtanx +tany=(m+n)tan1/2(x+y)

Grade:10

2 Answers

Aditya Gupta
5 years ago
lets put x=y=pi/4 and m=n=2
but mtanx +tany= m+1=3
and (m+n)tan1/2(x+y)=2tanpi/4=2
we see that 2 is not equal to3, so question is wrong
Aditya Gupta
5 years ago
 
lets put x=y=pi/4 and m=n=2
but mtanx +tany= m+1=3
and (m+n)tan1/2(x+y)=4tanpi/4=4
we see that 4 is not equal to3, so question is wrong